### v2

# Isotropic Covariance Functions for Disparate Spatial Resolutions

Abstract
Distances between sets arise naturally in data fusion problems involving both point-referenced and areal observations, as well as in set-indexed stochastic processes more broadly. However, commonly used constructions of distances on sets, including those derived from the Hausdorff distance, generally fail to be conditionally negative definite, precluding their use in isotropic covariance models. We propose the ball–Hausdorff distance, defined as the Hausdorff distance between the minimum enclosing balls of bounded sets in a metric space. For length spaces, we derive an explicit representation of this distance in terms of the associated centers and radii. We show that the ball–Hausdorff distance is conditionally negative definite whenever the underlying metric is conditionally negative definite, which implies, via Schoenberg’s theorem, an isometric embedding into a Hilbert space. As a consequence, broad classes of isotropic covariance functions, including the Matérn and powered exponential families, are valid for random fields indexed by sets. The resulting construction reduces set-to-set dependence to low-dimensional geometric summaries, leading to substantial computational simplifications in covariance evaluation.
Keywords: Change of support, Geostatistics, Hausdorff distance, Length spaces, Random fields, Spatial data fusion
1 Introduction
Many modern spatial analyses involve observations recorded over spatial units of differing locations, sizes, and geometries, creating fundamental challenges for modeling dependence across spatial supports. Such settings are common in environmental science and related areas such as air pollution epidemiology (e.g., Ellis-Soto et al.,, 2023; Zhong et al.,, 2024), where different aspects of the same underlying process are observed at multiple spatial resolutions. Several closely related statistical problems arise from this mismatch. Change of support refers to predicting a process indexed on one spatial scale using observations collected on another (Gelfand et al.,, 2001), with downscaling denoting the special case in which information aggregated over coarse spatial units is used to infer values at a finer resolution (Zheng et al.,, 2025). When observations on the same phenomenon are simultaneously available at multiple spatial resolutions, the problem is commonly framed as spatial data fusion (Moraga et al.,, 2017). In addition, response and explanatory variables are often observed on incompatible spatial supports, a setting known as spatial misalignment (Gotway and Young,, 2002). Despite differing terminology, these problems share a common structure: the need to model dependence between spatial units that differ in their spatial support.
Most existing approaches to multi-resolution spatial data model dependence by treating observations at different supports as aggregations of a common underlying continuous process, which leads to substantial computational and theoretical limitations. Areal measurements are typically represented as integrals of an underlying stochastic process (Fuentes and Raftery,, 2005). However, the implied finite-dimensional probability density functions involve stochastic integrals that are analytically intractable even when the underlying process is a Gaussian Process (GP). Consequently, inferences rely on numerical approximations (Cowles et al.,, 2002) or Monte Carlo integration (Gelfand et al.,, 2001), which can be computationally demanding and introduce hard-to-quantify bias (Gonçalves and Gamerman,, 2018). An alternative strategy avoids aggregation by defining stochastic processes directly on collections of spatial sets. The Hausdorff distance (Hausdorff,, 2005, pg. 167) has been explored as a basis for such constructions (Godoy et al.,, 2025), but its direct construction does not guarantee positive definite (PD) functions, leaving the existence and validity of the induced process unresolved. This lack of theoretical guarantee, coupled with high computational costs (Knauer et al.,, 2011), necessitates a more robust framework for set-to-set dependence.
Constructing valid PD functions becomes substantially more delicate once the underlying index space departs from Euclidean geometry. While many PD functions are known for $\mathbb{R}^{p}$ equipped with the Euclidean distance (Cressie,, 1993, Ch. 2.5.1), their positive definiteness may not extend to other spaces (Li et al.,, 2023). In fact, a function that is PD on $\mathbb{R}^{1}$ may not remain PD on $\mathbb{R}^{2}$, although the reverse is true (Cressie,, 1993, p. 22). This geometric sensitivity has motivated recent work on characterizing PD covariance functions on non-Euclidean spaces. The particular case of the sphere has been extensively studied (Banerjee,, 2005; Gneiting,, 2013; Guinness and Fuentes,, 2016; Porcu et al.,, 2021). Certain covariance functions exhibit the notable property of being PD on both the (p1)(p-1)-dimensional sphere ($\mathbb{S}^{p-1}$) and $p$-dimensional Euclidean space (Porcu et al.,, 2016). More recently, analogous questions have arisen for covariance functions indexed by graphs and their edges (Anderes et al.,, 2020). Despite these advances, there is no general construction of distances on spatial sets that both ensures positive definiteness of isotropic covariance functions and avoids the computational burdens associated with aggregation-based models.
To address this gap, we propose a framework for constructing isotropic covariance functions directly on collections of spatial sets, which accommodates point-referenced, areal, and mixed spatial data within a unified formulation. The goal is to circumvent the analytical and computational intractability of aggregation-based models while ensuring mathematical validity. To this end, we define a pseudometric, the ball-Hausdorff distance, based on the Hausdorff distance between minimum enclosing balls of sets. Under mild assumptions, we derive a closed-form expression for the proposed distance and show that it is conditionally negative definite (CND) whenever the underlying domain metric is CND. By Schoenberg’s theorem (Schoenberg,, 1938), this implies that the square-root of the ball-Hausdorff distance admits an isometric embedding into a Hilbert space, thereby providing a rich class of valid isotropic covariance functions indexed by spatial sets. As we demonstrate in numerical studies, this theoretical validity is a critical distinction, as the standard Hausdorff distance frequently yields invalid covariance structures in identical settings.
The remainder of the paper is organized as follows. Section 2 establishes the necessary preliminary results, including the properties of CND functions and their relationship with isometric embeddings on Hilbert spaces. Section 3 presents our main theoretical contribution: the definition of the ball-Hausdorff distance, the proof of its CND property for length spaces, and valid covariance functions for this distance. In Section 4, we provide numerical benchmarks across Euclidean and spherical domains to analyze the agreement between the ball-Hausdorff and standard Hausdorff distances, quantify computational speedups, and empirically show the positive definiteness of the resulting covariance matrices. Finally, we conclude with a discussion of the results and future directions for research in Section 5.
2 Preliminaries
Let $D$ denote an index set and $d\,:\,D\times D\to[0,\infty)$ a distance function. The pair $(D,d)$ is a pseudometric space if $d$ satisfies all the metric axioms except the identity of indiscernibles. That is, $d(x,y)=0$ does not necessarily imply that $x$ and $y$ are identical. In the literature, these functions are also termed semi-metrics (e.g., Wells and Williams,, 1975; Burago et al.,, 2001). We denote open and closed balls of radius $r$ by
[…]
respectively. We follow the convention of defining the distance between a point and a set, say $x\in D$ and $A\subset D$, as the infimum distance between them. Formally, $d(x,A)=\inf_{a\in A}d(x,a)$. Finally, the $r$-neighborhood of a set $A\subset D$ is the union of open balls centered at the elements of that set:
$\mathrm{B}_{r}(A)=\bigcup_{a\in A}\mathrm{B}_{r}(a).$
From here on, we focus on length spaces, where the distance is defined as the infimum of the lengths of admissible paths (i.e., continuous rectifiable curves) connecting two points (Burago et al.,, 2001, Ch. 2). This assumption is imposed to ensure the existence of midpoints (Burago et al.,, 2001, Lemma 2.4.8), a property that enables the derivation of the linear expansion of balls.
Lemma 2.1 (Linear expansion of balls).
Let $(D,d)$ be a length space. For any $x\in D$ and radii $r,k\geq 0$,
$\mathrm{B}_{r}\left(\mathrm{B}_{k}(x)\right)=\mathrm{B}_{r+k}(x).$
This result serves as a cornerstone for our main findings in Section 3. A proof is provided in Appendix A. For additional background on length spaces, we refer to Appendix B.
2.1 Conditionally negative definite functions and embeddings
Conditionally negative definite distance functions are closely related to PD isotropic covariance functions. A function $g\,:\,D\times D\to\mathbb{R}_{+}$ is CND, if it satisfies
$\sum_{i=1}^{m}\sum_{j=1}^{m}b_{i}b_{j}g(s_{i},s_{j})\leq 0,$ (1)
for $s_{1},\ldots,s_{m}\in D$ and any real numbers $b_{1},\ldots,b_{m}$ such that $\sum_{i}b_{i}=0$ (Rao,, 1984). In contrast, a PD function $r\,:\,D\times D\to\mathbb{R}$ must satisfy
$\sum_{i=1}^{m}\sum_{j=1}^{m}a_{i}a_{j}r(s_{i},s_{j})\geq 0,$
for any real numbers $a_{1},\ldots,a_{m}$. Finally, a real function $\varphi$ defined on non-negative real numbers is radial positive definite (RPD) on the pseudometric space $(D,d)$ provided that $\varphi$ is continuous and
$\sum_{i=1}^{m}\sum_{j=1}^{m}a_{i}a_{j}\varphi(d(s_{i},s_{j}))\geq 0,$ (2)
for real numbers $a_{1},\ldots,a_{m}$ and $s_{1},\ldots,s_{m}\in D$. Note that isotropic covariance functions represent a special case of RPD functions.
Denote $\Phi_{p}$ as the class RPD functions on $\mathbb{R}^{p}$. Functions belonging to this class satisfy (Schoenberg,, 1938; Gneiting,, 2013):
$\Phi_{1}\supset\Phi_{2}\supset\cdots\supset\Phi_{\infty}=\bigcap_{k=1}^{\infty}\Phi_{k}.$
Well known members of the class $\Phi_{\infty}$ include the Matérn, powered exponential (PEXP), generalized Cauchy, and Dagum families (Gneiting,, 2013, Table 2). For further properties and a detailed definition, we refer to Gneiting, (2013).
Isometric embeddings, a special class of mappings between spaces, are crucial to establish our results. An isometric embedding preserves the distances between points in the original space when mapped to the target space. More formally, a space $(D,d)$ can be isometrically embedded in another space $(D^{\ast},d^{\ast})$ if there exists a map $\phi\,:\,D\to D^{\ast}$ such that
$d^{\ast}(\phi(s),\phi(t))=d(s,t),$
for any $s,t\in D$ (Wells and Williams,, 1975).
The following theorem, originally due to Schoenberg, (1938), and later summarized by Wells and Williams, (1975), provides a fundamental link between conditional negative definiteness and Hilbert space embeddings. This result will play a central role in our construction of valid covariance functions on spaces indexed by sets.
Theorem 2.2.
A pseudometric space $(D,d)$ can be isometrically embedded in a Hilbert space if and only if $d^{2}$ is CND.
The central implication of this result for our purposes is that conditional negative definiteness directly enables the construction of valid isotropic covariance functions via Hilbert space embeddings. Proofs and further discussion of this classical result can be found in Wells and Williams, (1975, Theorem 2.4 and Remark 3.2) with additional review and commentary in Menegatto et al., (2020). The following corollary formalizes this connection for a widely used class of covariance models.
Corollary 2.2.1.
Let $(D,d)$ be a pseudometric space such that $d$ is CND. Then, $\exp\{-\kappa d^{\nu}\}$ is RPD for $\nu\in(0,1]$ and $\kappa>0$.
The proof is presented in Appendix A. Corollary 2.2.1 establishes that the PEXP function with power $\nu\in(0,1]$ is always positive definite on metric spaces equipped with CND metrics (on the original scale).
We now generalize the embedding-based argument to show that broad families of isotropic covariance functions remain valid on non-Euclidean spaces equipped with conditionally negative definite metrics. Specifically, Theorem 2.3 extends Theorem 2.2 to the class of radial positive definite functions and provides sufficient conditions for constructing valid covariance models beyond Euclidean domains.
Theorem 2.3.
Let $D$ be a non-Euclidean space and $d$ be a CND pseudometric. Then, any function belonging to the class $\Phi_{\infty}$ is RPD on $(D,d^{1/2})$.
Proof.
From Theorem 2.2, we have that $(D,d^{1/2})$ is a metric space that can be isometrically embedded in $H$. Therefore, functions in the class $\Phi_{\infty}$ are also RPD on $(D,d^{1/2})$. ∎
This result shows that covariance families in $\Phi_{\infty}$, originally defined for Hilbert spaces, remain valid on any domain equipped with a conditionally negative definite pseudometric after a square-root transformation of the distance. This implication allows covariance constructions developed in Euclidean settings to be transferred directly to a wide range of non-Euclidean domains encountered in spatial statistics, including those based on $\alpha$-norms (Curriero,, 2006) and graph-indexed spaces (Anderes et al.,, 2020).
Verifying whether a distance is conditionally negative definite can be challenging in general, but several established results provide useful guidance. For instance, any $\alpha$-norm on $\mathbb{R}^{2}$ is conditionally negative definite (Richards,, 1985). In higher dimensions, that is, $\mathbb{R}^{p}$ with $p\geq 3$, $\alpha$-norms are CND for $\alpha\in(0,2]$. In addition, Theorem 3.6 of (Meckes,, 2013) shows that if a metric $d$ is CND, then so is $d^{\nu}$ for $\nu\in(0,1]$, and that the great-circle distance on round spheres is also CND. In the construction developed in Section 3, however, conditional negative definiteness can be established directly from the structure of the proposed distance, substantially simplifying this verification step for set-indexed domains.
3 Isotropic positive definite functions for sets
To motivate the proposed distance, we first consider the classical Hausdorff distance, which provides a natural notion of discrepancy between spatial sets but lacks general guarantees for valid covariance construction. For non-empty sets $A_{1},A_{2}\subset D$, the Hausdorff distance is defined as
$d_{h}(A_{1},A_{2})=\inf\{r\geq 0\,:\,A_{1}\subseteq\mathrm{B}_{r}(A_{2}),A_{2}\subseteq\mathrm{B}_{r}(A_{1})\}.$ (3)
If no such $r$ exists, set $d_{h}(A_{1},A_{2}):=\infty$. While $d_{h}$ constitutes a valid metric on the class of non-empty, closed, and bounded subsets of $D$ (Arutyunov et al.,, 2016), it presents significant challenges for statistical modeling. Despite its intuitive interpretation in spatial statistics (Godoy et al.,, 2022, 2025), computing the Hausdorff distance is often prohibitively expensive (Knauer et al.,, 2011; Parikh,, 2023). More critically, to the best of our knowledge, constructing valid positive-definite functions based directly on $d_{h}$ remains an open problem. These analytical and computational limitations motivate the ball-Hausdorff distance defined below.
3.1 The ball-Hausdorff distance
We define the ball–Hausdorff distance by reducing each spatial set to its minimal enclosing ball, thereby retaining only location and scale information relevant for dependence modeling. Let $(D,d)$ be a pseudometric space and let $\mathcal{C}_{D}$ denote the class of bounded subsets of $D$. For any $A\in\mathcal{C}_{D}$ and $x\in D$, we define
${\rm R}(x,A)=\inf\{r\geq 0\,:\,A\subseteq\mathrm{B}_{r}(x)\}$
as the radius of the smallest ball centered at $x$ containing $A$.
To construct a unique minimum enclosing ball, we must rely on a notion of centrality. We adopt the Chebyshev center (Garkavi,, 1970, p. 128) because it directly minimizes the covering radius, ensuring the resulting summary is the smallest possible ball containing $A$.
Definition 3.1 (Chebyshev center).
The Chebyshev radius of any $A\in\mathcal{C}_{D}$, denoted by ${\rm R}(A)$, is the infimum of the radii of all balls in $D$ containing $A$:
${\rm R}(A)=\inf_{x\in D}{\rm R}(x,A).$
The set of Chebyshev centers of $A$, denoted by $\mathcal{E}(A)$, consists of the points in $D$ that achieve this minimum radius:
$\mathcal{E}(A)=\{x\in D:{\rm R}(x,A)={\rm R}(A)\}.$
If $\mathcal{E}(A)$ is non-empty, any element $c(A)\in\mathcal{E}(A)$ is referred to as a Chebyshev center of $A$.
We must address the existence and uniqueness of the Chebyshev center to ensure the metric is well-defined. Regarding existence, while the set of centers $\mathcal{E}(A)$ may be empty in arbitrary metric spaces, the Hopf–Rinow–Cohn-Vossen Theorem (Burago et al.,, 2001, Theorem 2.5.28) guarantees that in complete manifolds (such as the sphere and torus), closed and bounded sets are compact. Consequently, the radius function attains its minimum, ensuring $\mathcal{E}(A)$ is non-empty. Regarding uniqueness, the strict convexity of the norm in Euclidean spaces guarantees that $\mathcal{E}(A)$ is a singleton (Amir,, 1978). However, uniqueness is not guaranteed globally in general length spaces. For example, a great circle on $\mathbb{S}^{2}$ admits multiple centers (the poles). Therefore, to ensure validity in non-Euclidean settings, we impose further restrictions on $\mathcal{C}_{D}$. For the unit sphere, this requires a Chebyshev radius less than $\pi/2$ (Bridson and Haefliger,, 1999, Proposition 1.4). This condition satisfied by all spatial configurations in our numerical experiments.
Using the definition 3.1, we define the geometric summary $\mathcal{B}(A)$ as the minimum enclosing ball of $A$, characterized by the center $c(A)$ and radius $R(A)$. Note that, if $A$ is already a ball (or a singleton), it follows trivially that $A=\mathcal{B}(A)$.
Now we define the ball-Hausdorff distance between two bounded sets.
Definition 3.2.
The ball-Hausdorff distance between two sets $A_{1}$ and $A_{2}$ is defined as
$bh(A_{1},A_{2})=d_{h}(\mathcal{B}(A_{1}),\mathcal{B}(A_{2})),$
where $d_{h}(\cdot,\cdot)$ denotes the Hausdorff distance (see Equation (3)) and $\mathcal{B}(\cdot)$ denotes a minimum enclosing ball.
The ball–Hausdorff distance inherits basic pseudometric properties from the Hausdorff distance while remaining insensitive to within-set geometry. In particular, for $A_{1},A_{2},A_{3}\in\mathcal{C}_{D}$,
$bh(A_{1},A_{2})$ $\geq 0,$
$bh(A_{1},A_{2})$ $=0,\quad\text{ if }\quad A_{1}=A_{2},$
$bh(A_{1},A_{2})$ $\leq bh(A_{1},A_{3})+bh(A_{2},A_{3}),$
where the first two properties are immediate from the definition and the third follows from the fact that Hausdorff distance is a metric on $\mathcal{C}_{D}$ (Molchanov,, 2005, Appendix C). Note that the distance does not render a metric because $bh(A_{1},A_{2})=0$ does not imply $A_{1}$ and $A_{2}$ are identical. Importantly, the ball-Hausdorff distance reduces to the underlying metric $d$ when applied to singletons. In other words, the proposed distance function between points reduces to $d$, from the original metric space. In Appendix C, we also derive an upper-bound for the ball-Hausdorff distance.
An important question in practice is how to compute the ball-Hausdorff distance. At first sight, it involves computationally expensive operations. For instance, calculating the Hausdorff distance between two balls may not always be trivial (see Section 2 for the problems in calculating the Hausdorff distance). In the special case where the study region lies on the Euclidean plane, a closed-form formula for the Hausdorff distance between circles has been developed (Scitovski and Marošević,, 2015). We extend this formula to length spaces which is general enough to cover most scenarios relevant for spatial statistics applications.
Theorem 3.3.
Let $(D,d)$ be a length-space. Define $\mathcal{C}_{D}$ as the class of non-empty, bounded sets in $D$. For any $A_{1},A_{2}\in\mathcal{C}_{D}$, the ball-Hausdorff distance is:
$bh(A_{1},A_{2})=d(\rm{c}(A_{1}),\rm{c}(A_{2}))+\lvert{\rm R}(A_{1})-{\rm R}(A_{2})\rvert$.
Proof.
For ease of notation, let $x=\rm{c}(A_{1})$, $y=\rm{c}(A_{2})$, $r={\rm R}(A_{1})$ and $k={\rm R}(A_{2})$. Then,
$bh(A_{1},A_{2})=d_{h}(\mathrm{B}_{r}(x),\mathrm{B}_{k}(y)).$
By definition,
$d_{h}(\mathrm{B}_{r}(x),\mathrm{B}_{k}(y))$ $=\inf\{m\geq 0\,:\,\mathrm{B}_{r}(x)\subseteq\mathrm{B}_{m}(\mathrm{B}_{k}(y)),\mathrm{B}_{k}(y)\subseteq\mathrm{B}_{m}(\mathrm{B}_{r}(x))\}$
$=\inf\{m\geq 0\,:\,\mathrm{B}_{r}(x)\subseteq\mathrm{B}_{m+k}(y),\mathrm{B}_{k}(y)\subseteq\mathrm{B}_{m+r}(x)\},$
where the second line follows from Lemma 2.1.
The inclusion $\mathrm{B}_{r}(x)\subseteq\mathrm{B}_{k+m}(y)$ holds if and only if every point at distance less than $r$ from $x$ lies within distance $k+m$ from $y$. Formally,
$\sup_{u\in\mathrm{B}_{r}(x)}d(u,y)\leq k+m.$
Note that, for any $u\in\mathrm{B}_{r}(x)$,
$d(u,y)\leq d(u,x)+d(x,y)<r+d(x,y).$
Thus,
$r+d(x,y)\leq k+m\iff m\geq d(x,y)+r-k.$
Similarly, the inclusion $\mathrm{B}_{r}(x)\subseteq\mathrm{B}_{k+m}(y)$ holds if and only if
$m\geq d(x,y)+k-r.$
Thus, the smallest $m\geq 0$ satisfying both conditions is:
$m\geq\max\{d(x,y)+r-k,d(x,y)+k-r\}=d(x,y)+\lvert r-k\rvert$,
by the $\max$ formula. Therefore,
$bh(A_{1},A_{2})=d(\rm{c}(A_{1}),\rm{c}(A_{2}))+\lvert\rm{R}(A_{1})-\rm{R}(A_{2})\rvert$,
as claimed. ∎
Theorem 3.3 offers both a practical computational method and a clear geometric interpretation. Unlike the standard Hausdorff distance, which depends on extremal point-to-point discrepancies, the ball–Hausdorff distance depends only on relative location and scale through the sets’ Chebyshev centers and radii. In the limiting case where sets have identical radii or are small relative to their separation, the distance reduces to the distance between their centers.
3.2 Isotropic positive definite functions for sets
In addition to providing a simplified expression for the ball-Hausdorff distance, Theorem 3.3 enables a direct characterization of its conditional negative definiteness, which underpins our main result.
Theorem 3.4.
Let $(D,d)$ be a length-space where the distance $d$ is CND. Define $\mathcal{C}_{D}$ as the class of non-empty, bounded sets in $D$. Then,
  1. 1.
    $bh(\cdot,\cdot)$ is CND,
  2. 2.
    $(\mathcal{C}_{D},\sqrt{bh})$ can be embedded on a Hilbert space.
Proof.
Let us start with (1). From Theorem 3.3, we have
$bh(A_{1},A_{2})=d(\rm{c}(A_{1}),\rm{c}(A_{2}))+\lvert{\rm R}(A_{1})-{\rm R}(A_{2})\rvert$,
where $d({\rm c}(A_{1}),{\rm c}(A_{2}))$ is CND by assumption. The term $\lvert{\rm R}(A_{1})-{\rm R}(A_{2})\rvert$ stands for a $\alpha$-norm, with $\alpha=1$, on $\mathbb{R}^{1}$, which is also CND (Richards,, 1985). The family of CND functions form a convex cone (Cressie,, 1993, pg. 88). That is, linear combinations of CND functions that employ positive coefficients result in CND functions (Kapil et al.,, 2018). Thus, under the stated assumptions, $bh(A_{1},A_{2})$ is CND. The second part of the Theorem follows directly from Theorem 2.2. ∎
Combining Theorems 3.4 and 2.3 allows us to immediately establish positive definite functions for sets. In particular, for any length space equipped with a CND distance function, we can define valid positive definite functions. That follows from the fact that, in those cases, the square-root transformed ball-Hausdorff distance (on the class of bounded sets of these topological spaces) can be embedded on a Hilbert space. Thus, families of valid isotropic covariance functions on Hilbert spaces (i.e., functions belonging to the class $\Phi_{\infty}$) also render valid families for the proposed distance.
4 Numerical Examples
Building on the theoretical results in Section 3, we present numerical examples that validate the proposed ball–Hausdorff construction by comparing it with the Hausdorff distance, assessing computational efficiency, and empirically verifying positive definiteness of the resulting covariance matrices. To this end, we consider two widely adopted isotropic covariance families: the PEXP family,
$C(h;\phi,\nu)=\exp\left\{-\frac{h^{\nu}}{\phi^{\nu}}\right\},$
and the Matérn family,
$C(h;\phi,\nu)=\frac{2^{\nu-1}}{\Gamma(\nu)}{\left(\frac{h}{\phi}\right)}^{\nu}K_{\nu}\left(\frac{h}{\phi}\right),$
where $h$ represents the distance between two elements from the index set, $\phi$ controls the strength of the spatial dependence, and $\nu$ governs smoothness. Both families belong to the class $\Phi_{\infty}$ and are therefore valid on Hilbert spaces. For interpretability, dependence is summarized using the practical range $\rho$, defined as the distance where correlation decays to 0.1. For the PEXP family, $\rho={\log(10)}^{1/\nu}\phi$ (Godoy et al.,, 2025), while for the Matérn family, $\rho\approx\phi\sqrt{8\nu}$ (Lindgren et al.,, 2011). These models form the basis for distance comparisons and eigenvalue-based diagnostics that follow.
Figure 1: Synthetic spatial configurations used in the numerical examples. Rows correspond to the Euclidean plane $(\mathbb{R}^{2})$ and the sphere $(\mathbb{S}^{2})$, while columns represent gridded, fused, and areal settings. These examples are designed to span increasing spatial complexity across domains commonly encountered in multi-resolution spatial analyses.
To empirically evaluate the proposed framework, we consider synthetic datasets designed to reflect three common spatial configurations encountered in practice: gridded data, typical of satellite-derived products; fused data, combining point-referenced and areal observations; and areal data consisting of irregular polygons. These configurations are constructed on both the Euclidean plane $(\mathbb{R}^{2})$ and the sphere $(\mathbb{S}^{2})$ to assess performance across distinct geometric domains (Figure 1). Using these settings, our analysis pursues three objectives: first, to assess empirical agreement between the proposed ball-Hausdorff distance and the standard Hausdorff distance; second, to benchmark computational efficiency arising from the closed-form representation of the proposed distance; and third, to verify positive definiteness of covariance matrices constructed under the ball-Hausdorff metric, in contrast to those based on the Hausdorff distance.
Figure 2: wo-dimensional kernel density estimates comparing pairwise Hausdorff distances (x-axis) and ball–Hausdorff distances (y-axis) across spatial configurations and domains. Columns correspond to gridded, fused, and areal settings, while rows represent the Euclidean plane and the sphere. Distances are rescaled to the unit interval within each scenario. Concentration of mass along the dashed $y=x$ line indicates close empirical agreement between the two distance measures across increasing spatial complexity.
We computed pairwise distance matrices for all configurations shown in Figure 1 using both the ball-Hausdorff and Hausdorff distances. Distances were rescaled to the unit interval by dividing by the maximum observed distance within each scenario, so that comparisons reflect relative rather than absolute scale differences across spatial configurations. Under this joint normalization, the values for of either metric need not span the full unit interval, as the observed minimum and maximum can be bounded away from 0 and 1 by geometric constraints inherent to the synthetic designs, such as shared boundaries, partial overlap, and limited variability in set geometry. Figure 2 displays two-dimensional kernel density estimates comparing the resulting distance values. The concentration of mass along the $y=x$ diagonal indicates close empirical agreement between the two distances, with no systematic curvature or dispersion as spatial complexity increases from gridded to fused and irregular areal units. In particular, the near-monotone relationship across all configurations suggests that the ball-Hausdorff distance preserves the relative proximity structure induced by the Hausdorff distance, despite relying on a substantially reduced geometric representation.
The two distances differ substantially in computational cost. Evaluation of the Hausdorff distance requires identifying extremal point-to-point discrepancies between spatial sets, which becomes increasingly expensive as set complexity and resolution increase. In contrast, the ball-Hausdorff distance admits a closed-form expression that depends only on Chebyshev centers and radii (Theorem 3.3), eliminating the need for pairwise pointwise comparisons. This structural reduction leads to substantial computational savings. In our experiments, evaluation based on the ball-Hausdorff distance was between 11 and 183 times faster than evaluation based on the Hausdorff distance, with the largest speedups observed for areal data on the sphere.
We conclude the numerical examples with an eigenvalue-based comparison of covariance matrices constructed using the ball-Hausdorff and standard Hausdorff distances. Two covariance families are considered: the PEXP family with $\nu=1$ and the Matérn family with $\nu=2.5$. For the Matérn family, covariance matrices are constructed using square-root transformed distances, as required by Theorem 2.3. The practical range parameter is varied over $\rho\in[0.8,1.5]$, where the difference between the positive definiteness of the resulting covariance matrices from the two distances are notable. The smallest eigenvalue of each resulting covariance matrix, denoted $\lambda_{1}$, is used to assess positive definiteness, with negative values indicating invalid covariance structures.
Figure 3: Eigenvalue diagnostics for covariance matrices constructed using the PEXP family with $\nu=1$ (top row) and the Matérn family with $\nu=2.5$ (bottom row). For each panel, the smallest eigenvalue $\lambda_{1}$ is plotted as a function of the practical range $\rho$ across spatial configurations (columns) and domains (rows). Solid red lines correspond to covariance matrices based on the ball-Hausdorff distance, while dashed blue lines correspond to those based on the Hausdorff distance. For the Matérn family, square-root transformed distances are used as required by Theorem 2.3. The gray shaded region indicates invalid (non-positive definite) covariance structures ($\lambda_{1}<0$).
Figure 3 summarizes the resulting eigenvalue diagnostics across spatial configurations and domains. In all cases considered, covariance matrices based on the ball-Hausdorff distance remain positive definite, with $\lambda_{1}\geq 0$ throughout the examined range of $\rho$. In contrast, covariance matrices constructed using the standard Hausdorff distance frequently exhibit negative eigenvalues, even when square-root transformation is applied for the Matérn family, a reported limitations of Hausdorff-based constructions in set-indexed settings (Godoy et al.,, 2025). These violations are most pronounced for fused data and for configurations defined on the sphere. Collectively, these results serve as a final numerical confirmation that replacing the Hausdorff distance with the ball-Hausdorff distance leads to systematically improved validity of covariance constructions in set-indexed spatial settings.
5 Discussion
We introduced the ball-Hausdorff distance, a new distance function for spatial sets, and established its theoretical properties in set-indexed settings. Under the assumption that the underlying domain is a length space, the proposed distance admits a closed-form expression that depends only on the Chebyshev centers and radii of the sets. This structural reduction allows conditional negative definiteness of the base metric to be inherited by the ball-Hausdorff distance, which in turn guarantees the validity of broad classes of isotropic covariance functions through Hilbert-space embeddings. Our numerical results demonstrate that, although the classic Hausdorff distance provides an intuitive notion of discrepancy between sets, it frequently fails to yield valid covariance matrices, particularly on non-Euclidean domains such as the sphere. In contrast, the ball-Hausdorff distance ensures theoretical validity while remaining compatible with commonly used covariance families.
The closed-form expression induced by the geometric reduction of the ball–Hausdorff distance leads directly to substantial computational simplification. By representing each set through its center and radius, evaluation of the distance avoids the extremal point-to-point calculations required by the Hausdorff distance. The resulting computational gains observed in our numerical studies, including speedups of over 180 times in areal data configurations, follow from this structural simplification rather than from approximation or numerical shortcuts. As a consequence, the proposed framework remains applicable in large-scale settings involving many spatial units or high-resolution partitions, enabling covariance modeling over large collections of spatial sets without the need for computationally intensive distance calculations or numerical evaluation of stochastic integrals.
The application of covariance families under the proposed framework depends on the scale at which the ball–Hausdorff distance is transformed to preserve positive definiteness. While the PEXP family with $\nu\in(0,1]$ can be applied directly to the ball-Hausdorff distance, other families, such as the Matérn class, require the square-root transformed distance to ensure theoretical validity (Theorem 2.3). Because the square-root is a monotone transformation, inference on range-related parameters remains identifiable on the original distance scale. In frequentist settings, invariance of the maximum likelihood estimator permits reporting the practical range on the original scale, with uncertainty quantified, for example, via the delta method. In Bayesian analyses, the same transformation may be applied to posterior samples or handled explicitly through a Jacobian adjustment. These properties allow direct construction of valid set-indexed Gaussian processes for problems such as spatial data fusion and change-of-support, while still accommodating stochastic integrals when scientifically appropriate. An interesting theoretical direction is the development of valid cross-covariance functions for sets (Genton and Kleiber,, 2015), which would support modeling when responses and explanatory variables are observed on different spatial supports.
Several theoretical questions remain open. One concerns the mean-square differentiability of random fields defined through the proposed covariance construction. For the standard Matérn family, differentiability is governed by the smoothness parameter (Stein,, 1999, pg. 32), but extending this notion to set-indexed or manifold-indexed processes requires more general concepts of differentiation (Khastan et al.,, 2021). This issue is non-trivial, even for Gaussian processes on spheres (Alegría et al.,, 2021; Guinness and Fuentes,, 2016), and while recent work has defined mean-square differentiability for Gaussian processes on manifolds using smooth paths (Lin et al.,, 2019), a systematic treatment in the present setting remains to be developed. A further limitation arises from the pseudometric nature of the ball–Hausdorff distance: distinct sets with identical centers and radii are indistinguishable, which may lead to singular covariance matrices in degenerate cases. As in many reduced-representation models, this issue can be addressed by introducing a nugget effect, a standard regularization in geostatistical modeling that accounts for measurement error and unresolved small-scale variability (Cressie,, 1993, pg. 59–60).
Acknowledgments
Fernando A. Quintana was partially funded by grant ANID Fondecyt Regular 1220017. Marcos O. Prates acknowledges the research grants obtained from CNPq-Brazil (309186/2021-8), FAPEMIG (APQ-01837-22, APQ-01748-24), and CAPES, respectively, for partial financial support.
Appendix A Proofs
A.1 Lemma 2.1
Proof.
Let $u\in\mathrm{B}_{r}(\mathrm{B}_{k}(x))$. Then $d(u,v)<r$ for some $v\in\mathrm{B}_{k}(x)$, so $d(x,v)<k$. By the triangle inequality,
$d(x,u)\leq d(x,v)+d(v,u)<k+r,$
thus, $u\in\mathrm{B}_{r+k}(x)$.
Conversely, let $u\in\mathrm{B}_{r+k}(x)$. Note that $(D,d)$ is a length space (by assumption). In length spaces shortest paths always exist (Burago et al.,, 2001, Lemma 2.4.8). Thus, let $\gamma$ be a shortest path from $x$ to $u$. Let $v$ be the point on $\gamma$ at distance $k$ from $x$. Then $d(x,v)=k$ and d(u,v)=d(x,u)k<rd(u,v)=d(x,u)-k<r. Therefore, $v\in\mathrm{B}_{k}(x)$ and $u\in\mathrm{B}_{r}(v)$, proving $\mathrm{B}_{r+k}(x)\subseteq\mathrm{B}_{r}(\mathrm{B}_{k}(x))$. ∎
A.2 Corollary 2.2.1
Proof.
If $d$ is CND, Theorem 2.2 ensures that $(D,d^{1/2})$ can be isometrically embedded in a Hilbert space $H$. Consequently, there exists a map $\phi:D\to\mathbb{R}^{p}$ such that $d(s,t)=\lVert\phi(s)-\phi(t)\rVert_{2}$ for all $s,t\in D$. By Corollary 4.8 of Wells and Williams, (1975), the function $\exp\{-\kappa\lVert\cdot\rVert_{2}^{2\nu}\}$ is RPD on $L_{2}$ for $\nu\in(0,1]$. Therefore, for any set of points $s_{1},\dots,s_{m}\in D$ and real numbers $a_{1},\dots,a_{m}$, we have
$\sum_{i=1}^{m}\sum_{j=1}^{m}a_{i}a_{j}\exp\{-\kappa d(s_{i},s_{j})^{\nu}\}$ $=\sum_{i=1}^{m}\sum_{j=1}^{m}a_{i}a_{j}\exp\{-\kappa\lVert\phi(s_{i})-\phi(s_{j})\rVert_{2}^{2\nu}\}\geq 0.$
This establishes that $\exp\{-\kappa d^{\nu}\}$ is RPD on $(D,d)$. ∎
Appendix B Length and geodesic spaces
A particularly important class of metric spaces are length spaces, where the distance is determined by paths connecting points. A path $\gamma$ in a space $(D,d)$ is a continuous map $\gamma\,:\,I\to D$ defined on any connected subset of the real line, that is, an interval $I\subset\mathbb{R}$. A length structure consists of a class of “admissible paths” (e.g., all continuous paths) for which we can measure a non-negative length.
Definition B.1 (Definition 2.3.1 in Burago et al., (2001)).
Let $(D,d)$ be a metric space and $\gamma$ a continuous map $\gamma\,:\,[a,b]\to D$. Consider a partition $Y=\{y_{0},\ldots,y_{N}\}$ of $[a,b]$ such that $a=y_{0}\leq y_{1}\leq y_{2}\leq\cdot\leq y_{N}=b$. Then, the length of $\gamma$ (with respect to $d$) is
$L_{d}(\gamma)=\sup_{Y}\Sigma(Y),$
where
$\Sigma(Y)=\sum_{i=1}^{N}d(\gamma(y_{i-1}),\gamma(y_{i})).$
Note that, the length of a path is additive and for further properties, we refer to Burago et al., (2001). Now, we may formalize a length space as follows. In our context, a length space is a metric space such that the associated notion of distance is given by the infimum of lengths of all admissible paths between any two elements $x,y\in D$. That is,
$d_{L}(x,y)=\inf\{L_{d}(\gamma)\,;\,\gamma:[a,b]\to D,\,\gamma\in A,\gamma(a)=x,\gamma(b)=y\}.$
When the construction above induces the initial metric $d$ (i.e., $d_{L}=d$), than it is said to be intrinsic (Burago et al.,, 2001).
Although our primary motivation for working with length spaces is theoretical, they are broad enough to encompass the practical statistical challenges we address. Prominent examples include all normed vector spaces (Proposition 1.6 from Bridson and Haefliger,, 1999), piecewise smooth paths on manifolds, and broken lines in $\mathbb{R}^{n}$ (Burago et al.,, 2001, pg 27). A special case is the geodesic space, where the distance infimum is always attained by a specific path, known as a geodesic or shortest path (Bridson and Haefliger,, 1999, Ch. I.3). While every geodesic space is a length space, the converse is not always true (Bridson and Haefliger,, 1999, Ch. I.3). It is also worth noting that shortest paths need not be unique.
Appendix C An upper bound for the ball-Hausdorff distance
In this section, we utilize an alternative formulation of the Hausdorff distance. Specifically, contrasting with Equation (3), we express the metric as (Arutyunov et al.,, 2016; Barrera et al.,, 2021)
$d_{h}(A_{1},A_{2})=\max\left\{\sup_{x\in A_{1}}\inf_{y\in A_{2}}d(x,y),\sup_{x\in A_{2}}\inf_{y\in A_{1}}d(x,y)\right\}.$ (4)
A closed-form upper bound for the ball-Hausdorff distance is available under the mild assumption that $d$ is a pseudometric.
Lemma C.1.
Suppose $(D,d)$ is a pseudometric space. Define $\mathcal{C}_{D}$ as the class of non-empty, bounded sets in $D$. Let $A_{i}\in\mathcal{C}_{D}$, with $i=1,2$. Then, an upper-bound for the ball-Hausdorff distance is as follows
$bh(A_{1},A_{2})\leq d({\rm c}(A_{1}),{\rm c}(A_{2}))+\max\{{\rm R}(A_{1}),{\rm R}(A_{2})\}.$
Proof.
Let $\mathcal{B}(A_{1})=B_{r_{1}}(c_{1})$ and $\mathcal{B}(A_{2})=B_{r_{2}}(c_{2})$, where $c_{i}=c(A_{i})$ and $r_{i}=R(A_{i})$ denote the unique centers and Chebyshev radii, respectively. By definition, $bh(A_{1},A_{2})=d_{h}(B_{r_{1}}(c_{1}),B_{r_{2}}(c_{2}))$.
First, consider the directed distance from $B_{r_{1}}(c_{1})$ to $B_{r_{2}}(c_{2})$. For any point $x\in B_{r_{1}}(c_{1})$, we have $d(x,c_{1})\leq r_{1}$. The distance from $x$ to the set $B_{r_{2}}(c_{2})$ satisfies:
$d(x,B_{r_{2}}(c_{2}))=\inf_{y\in B_{r_{2}}(c_{2})}d(x,y).$ (5)
Since the center $c_{2}$ is contained in $B_{r_{2}}(c_{2})$, it follows that:
$d(x,B_{r_{2}}(c_{2}))\leq d(x,c_{2}).$ (6)
Applying the triangle inequality:
$d(x,c_{2})\leq d(x,c_{1})+d(c_{1},c_{2})\leq r_{1}+d(c_{1},c_{2}).$ (7)
Taking the supremum over all $x\in B_{r_{1}}(c_{1})$:
$\sup_{x\in B_{r_{1}}(c_{1})}d(x,B_{r_{2}}(c_{2}))\leq d(c_{1},c_{2})+r_{1}.$ (8)
By symmetry, the directed distance from $B_{r_{2}}(c_{2})$ to $B_{r_{1}}(c_{1})$ satisfies:
$\sup_{y\in B_{r_{2}}(c_{2})}d(y,B_{r_{1}}(c_{1}))\leq d(c_{1},c_{2})+r_{2}.$ (9)
Combining these results into the definition of the Hausdorff distance:
$bh(A_{1},A_{2})$ $=\max\left\{\sup_{x\in B_{r_{1}}(c_{1})}d(x,B_{r_{2}}(c_{2})),\sup_{y\in B_{r_{2}}(c_{2})}d(y,B_{r_{1}}(c_{1}))\right\}$
$\leq\max\{d(c_{1},c_{2})+r_{1},d(c_{1},c_{2})+r_{2}\}$
$=d(c_{1},c_{2})+\max\{r_{1},r_{2}\}$
$=d(c(A_{1}),c(A_{2}))+\max\{R(A_{1}),R(A_{2})\}.$
This completes the proof. ∎

### v3

@@
-Isotropic Covariance Functions for Disparate Spatial Resolutions
+On the Validity of Isotropic Covariance Functions for Set-indexed Random Fields
 Abstract
-Distances between sets arise naturally in data fusion problems involving both point-referenced and areal observations, as well as in set-indexed stochastic processes more broadly.
+Distances between sets arise naturally when modeling stochastic dependence on collections of spatial supports, including settings with point-referenced and areal observations.
 However, commonly used constructions of distances on sets, including those derived from the Hausdorff distance, generally fail to be conditionally negative definite, precluding their use in isotropic covariance models.
@@
 For length spaces, we derive an explicit representation of this distance in terms of the associated centers and radii.
-We show that the ball–Hausdorff distance is conditionally negative definite whenever the underlying metric is conditionally negative definite, which implies, via Schoenberg’s theorem, an isometric embedding into a Hilbert space.
-As a consequence, broad classes of isotropic covariance functions, including the Matérn and powered exponential families, are valid for random fields indexed by sets.
-The resulting construction reduces set-to-set dependence to low-dimensional geometric summaries, leading to substantial computational simplifications in covariance evaluation.
-Keywords: Change of support, Geostatistics, Hausdorff distance, Length spaces, Random fields, Spatial data fusion
+We show that the ball–Hausdorff distance is conditionally negative definite whenever the underlying metric is conditionally negative definite.
+By Schoenberg’s theorem, this implies an isometric embedding into a Hilbert space and guarantees the validity of broad classes of isotropic covariance functions, including the Matérn and powered exponential families, for set-indexed random fields.
+The construction reduces dependence between sets to low-dimensional geometric summaries, leading to substantial simplifications in covariance evaluation.
+Keywords: Change of support, Geostatistics, Hausdorff distance, Length spaces, Spatial data fusion
 1 Introduction
-Many modern spatial analyses involve observations recorded over spatial units of differing locations, sizes, and geometries, creating fundamental challenges for modeling dependence across spatial supports.
-Such settings are common in environmental science and related areas such as air pollution epidemiology (e.g., Ellis-Soto et al.,, 2023; Zhong et al.,, 2024), where different aspects of the same underlying process are observed at multiple spatial resolutions.
+Many modern spatial analyses require defining dependence across observations indexed by spatial units with heterogeneous locations, sizes, and geometries.
 Several closely related statistical problems arise from this mismatch.
@@
 When observations on the same phenomenon are simultaneously available at multiple spatial resolutions, the problem is commonly framed as spatial data fusion (Moraga et al.,, 2017).
-In addition, response and explanatory variables are often observed on incompatible spatial supports, a setting known as spatial misalignment (Gotway and Young,, 2002).
-Despite differing terminology, these problems share a common structure: the need to model dependence between spatial units that differ in their spatial support.
+In addition, response and explanatory variables are often observed on incompatible spatial supports, a setting known as spatial misalignment (Godoy et al., 2026b, ).
+Although these settings are often treated separately, they expose a common methodological challenge: constructing valid dependence structures on collections of spatial sets rather than on points.
 Most existing approaches to multi-resolution spatial data model dependence by treating observations at different supports as aggregations of a common underlying continuous process, which leads to substantial computational and theoretical limitations.
@@
 However, the implied finite-dimensional probability density functions involve stochastic integrals that are analytically intractable even when the underlying process is a Gaussian Process (GP).
-Consequently, inferences rely on numerical approximations (Cowles et al.,, 2002) or Monte Carlo integration (Gelfand et al.,, 2001), which can be computationally demanding and introduce hard-to-quantify bias (Gonçalves and Gamerman,, 2018).
+Consequently, inferences rely on numerical approximations or Monte Carlo integration (Gelfand et al.,, 2001), which are computationally demanding even for moderately sized datasets and introduce hard-to-quantify bias (Gonçalves and Gamerman,, 2018).
 An alternative strategy avoids aggregation by defining stochastic processes directly on collections of spatial sets.
-The Hausdorff distance (Hausdorff,, 2005, pg.
-167) has been explored as a basis for such constructions (Godoy et al.,, 2025), but its direct construction does not guarantee positive definite (PD) functions, leaving the existence and validity of the induced process unresolved.
+The Hausdorff distance has been explored as a basis for such constructions (Godoy et al., 2026a, ), but its direct construction does not guarantee positive definite (PD) functions, leaving the existence and validity of the induced process unresolved.
 This lack of theoretical guarantee, coupled with high computational costs (Knauer et al.,, 2011), necessitates a more robust framework for set-to-set dependence.
@@
 This geometric sensitivity has motivated recent work on characterizing PD covariance functions on non-Euclidean spaces.
-The particular case of the sphere has been extensively studied (Banerjee,, 2005; Gneiting,, 2013; Guinness and Fuentes,, 2016; Porcu et al.,, 2021).
+The particular case of the sphere has been extensively studied (see Porcu et al., (2021) and references therein).
 Certain covariance functions exhibit the notable property of being PD on both the (p1)(p-1)-dimensional sphere ($\mathbb{S}^{p-1}$) and $p$-dimensional Euclidean space (Porcu et al.,, 2016).
@@
 Despite these advances, there is no general construction of distances on spatial sets that both ensures positive definiteness of isotropic covariance functions and avoids the computational burdens associated with aggregation-based models.
-To address this gap, we propose a framework for constructing isotropic covariance functions directly on collections of spatial sets, which accommodates point-referenced, areal, and mixed spatial data within a unified formulation.
-The goal is to circumvent the analytical and computational intractability of aggregation-based models while ensuring mathematical validity.
-To this end, we define a pseudometric, the ball-Hausdorff distance, based on the Hausdorff distance between minimum enclosing balls of sets.
-Under mild assumptions, we derive a closed-form expression for the proposed distance and show that it is conditionally negative definite (CND) whenever the underlying domain metric is CND.
+To address this gap, we propose a framework for constructing isotropic covariance functions directly on collections of spatial sets, accommodating point-referenced, areal, and mixed spatial data within a unified formulation.
+We aim to circumvent the analytical and computational intractability of aggregation-based models while ensuring mathematical validity.
+To this end, we define a pseudometric, the ball-Hausdorff distance, based on the Hausdorff distance between the minimum enclosing balls of sets.
+Under mild assumptions, we derive a closed-form expression for this distance, allowing for fast and exact computation.
+Moreover, the proposed distance is conditionally negative definite (CND) whenever the underlying domain metric is CND.
 By Schoenberg’s theorem (Schoenberg,, 1938), this implies that the square-root of the ball-Hausdorff distance admits an isometric embedding into a Hilbert space, thereby providing a rich class of valid isotropic covariance functions indexed by spatial sets.
@@
 In the literature, these functions are also termed semi-metrics (e.g., Wells and Williams,, 1975; Burago et al.,, 2001).
-We denote open and closed balls of radius $r$ by
-[…]
-respectively.
+We denote open balls of radius $r$ by $\mathrm{B}_{r}(x)=\{u\in D:d(x,u)<r\}$.
 We follow the convention of defining the distance between a point and a set, say $x\in D$ and $A\subset D$, as the infimum distance between them.
 Formally, $d(x,A)=\inf_{a\in A}d(x,a)$.
-Finally, the $r$-neighborhood of a set $A\subset D$ is the union of open balls centered at the elements of that set:
-$\mathrm{B}_{r}(A)=\bigcup_{a\in A}\mathrm{B}_{r}(a).$
+Finally, the $r$-neighborhood of a set $A\subset D$ is the union of open balls centered at the elements of that set: $\mathrm{B}_{r}(A)=\bigcup_{a\in A}\mathrm{B}_{r}(a)$.
 From here on, we focus on length spaces, where the distance is defined as the infimum of the lengths of admissible paths (i.e., continuous rectifiable curves) connecting two points (Burago et al.,, 2001, Ch.
@@
 $\mathrm{B}_{r}\left(\mathrm{B}_{k}(x)\right)=\mathrm{B}_{r+k}(x).$
-This result serves as a cornerstone for our main findings in Section 3. A proof is provided in Appendix A. For additional background on length spaces, we refer to Appendix B.
+Proof.
+Let $u\in\mathrm{B}_{r}(\mathrm{B}_{k}(x))$. Then $d(u,v)<r$ for some $v\in\mathrm{B}_{k}(x)$, so $d(x,v)<k$. By the triangle inequality,
+$d(x,u)\leq d(x,v)+d(v,u)<k+r,$
+thus, $u\in\mathrm{B}_{r+k}(x)$.
+Conversely, let $u\in\mathrm{B}_{r+k}(x)$.
+Note that $(D,d)$ is a length space (by assumption).
+In length spaces shortest paths always exist (Burago et al.,, 2001, Lemma 2.4.8).
+Thus, let $\gamma$ be a shortest path from $x$ to $u$.
+Let $v$ be the point on $\gamma$ at distance $k$ from $x$.
+Then $d(x,v)=k$ and d(u,v)=d(x,u)k<rd(u,v)=d(x,u)-k<r.
+Therefore, $v\in\mathrm{B}_{k}(x)$ and $u\in\mathrm{B}_{r}(v)$, proving $\mathrm{B}_{r+k}(x)\subseteq\mathrm{B}_{r}(\mathrm{B}_{k}(x))$.
+∎
+This result serves as a cornerstone for our main findings in Section 3. For additional background on length spaces, we refer to Appendix A.
 2.1 Conditionally negative definite functions and embeddings
 Conditionally negative definite distance functions are closely related to PD isotropic covariance functions. A function $g\,:\,D\times D\to\mathbb{R}_{+}$ is CND, if it satisfies
-$\sum_{i=1}^{m}\sum_{j=1}^{m}b_{i}b_{j}g(s_{i},s_{j})\leq 0,$ (1)
-for $s_{1},\ldots,s_{m}\in D$ and any real numbers $b_{1},\ldots,b_{m}$ such that $\sum_{i}b_{i}=0$ (Rao,, 1984). In contrast, a PD function $r\,:\,D\times D\to\mathbb{R}$ must satisfy
-$\sum_{i=1}^{m}\sum_{j=1}^{m}a_{i}a_{j}r(s_{i},s_{j})\geq 0,$
-for any real numbers $a_{1},\ldots,a_{m}$.
-Finally, a real function $\varphi$ defined on non-negative real numbers is radial positive definite (RPD) on the pseudometric space $(D,d)$ provided that $\varphi$ is continuous and
-$\sum_{i=1}^{m}\sum_{j=1}^{m}a_{i}a_{j}\varphi(d(s_{i},s_{j}))\geq 0,$ (2)
+$\sum_{i=1}^{m}\sum_{j=1}^{m}b_{i}b_{j}g(s_{i},s_{j})\leq 0,$
+for $s_{1},\ldots,s_{m}\in D$ and any real numbers $b_{1},\ldots,b_{m}$ such that $\sum_{i}b_{i}=0$.
+In contrast, a real function $\varphi$ defined on non-negative real numbers is radial positive definite (RPD) on the pseudometric space $(D,d)$ provided that $\varphi$ is continuous and
+$\sum_{i=1}^{m}\sum_{j=1}^{m}a_{i}a_{j}\varphi(d(s_{i},s_{j}))\geq 0,$
 for real numbers $a_{1},\ldots,a_{m}$ and $s_{1},\ldots,s_{m}\in D$. Note that isotropic covariance functions represent a special case of RPD functions.
-Denote $\Phi_{p}$ as the class RPD functions on $\mathbb{R}^{p}$. Functions belonging to this class satisfy (Schoenberg,, 1938; Gneiting,, 2013):
+Denote $\Phi_{p}$ as the class of RPD functions on $\mathbb{R}^{p}$. Functions belonging to this class satisfy (Schoenberg,, 1938; Gneiting,, 2013):
 $\Phi_{1}\supset\Phi_{2}\supset\cdots\supset\Phi_{\infty}=\bigcap_{k=1}^{\infty}\Phi_{k}.$
@@
 An isometric embedding preserves the distances between points in the original space when mapped to the target space.
-More formally, a space $(D,d)$ can be isometrically embedded in another space $(D^{\ast},d^{\ast})$ if there exists a map $\phi\,:\,D\to D^{\ast}$ such that
+More formally, a space $(D,d)$ can be isometrically embedded on another space $(D^{\ast},d^{\ast})$ if there exists a map $\phi\,:\,D\to D^{\ast}$ such that
 $d^{\ast}(\phi(s),\phi(t))=d(s,t),$
@@
 Theorem 2.2.
-A pseudometric space $(D,d)$ can be isometrically embedded in a Hilbert space if and only if $d^{2}$ is CND.
+A pseudometric space $(D,d)$ can be isometrically embedded on a Hilbert space if and only if $d^{2}$ is CND.
 The central implication of this result for our purposes is that conditional negative definiteness directly enables the construction of valid isotropic covariance functions via Hilbert space embeddings.
@@
 Let $(D,d)$ be a pseudometric space such that $d$ is CND. Then, $\exp\{-\kappa d^{\nu}\}$ is RPD for $\nu\in(0,1]$ and $\kappa>0$.
-The proof is presented in Appendix A.
+Proof.
+If $d$ is CND, Theorem 2.2 ensures that $(D,d^{1/2})$ can be isometrically embedded in a Hilbert space $H$.
+Consequently, there exists a map $\phi:D\to\mathbb{R}^{p}$ such that $d(s,t)=\lVert\phi(s)-\phi(t)\rVert_{2}$ for all $s,t\in D$.
+By Corollary 4.8 of Wells and Williams, (1975), the function $\exp\{-\kappa\lVert\cdot\rVert_{2}^{2\nu}\}$ is RPD on $L_{2}$ for $\nu\in(0,1]$.
+Therefore, for any set of points $s_{1},\dots,s_{m}\in D$ and real numbers $a_{1},\dots,a_{m}$, we have
+$\sum_{i=1}^{m}\sum_{j=1}^{m}a_{i}a_{j}\exp\{-\kappa{d(s_{i},s_{j})}^{\nu}\}$ $=\sum_{i=1}^{m}\sum_{j=1}^{m}a_{i}a_{j}\exp\{-\kappa\lVert\phi(s_{i})-\phi(s_{j})\rVert_{2}^{2\nu}\}\geq 0.$
+This establishes that $\exp\{-\kappa d^{\nu}\}$ is RPD on $(D,d)$, for $\nu\in(0,1]$. ∎
 Corollary 2.2.1 establishes that the PEXP function with power $\nu\in(0,1]$ is always positive definite on metric spaces equipped with CND metrics (on the original scale).
@@
 This result shows that covariance families in $\Phi_{\infty}$, originally defined for Hilbert spaces, remain valid on any domain equipped with a conditionally negative definite pseudometric after a square-root transformation of the distance.
-This implication allows covariance constructions developed in Euclidean settings to be transferred directly to a wide range of non-Euclidean domains encountered in spatial statistics, including those based on $\alpha$-norms (Curriero,, 2006) and graph-indexed spaces (Anderes et al.,, 2020).
-Verifying whether a distance is conditionally negative definite can be challenging in general, but several established results provide useful guidance.
-For instance, any $\alpha$-norm on $\mathbb{R}^{2}$ is conditionally negative definite (Richards,, 1985).
-In higher dimensions, that is, $\mathbb{R}^{p}$ with $p\geq 3$, $\alpha$-norms are CND for $\alpha\in(0,2]$.
-In addition, Theorem 3.6 of (Meckes,, 2013) shows that if a metric $d$ is CND, then so is $d^{\nu}$ for $\nu\in(0,1]$, and that the great-circle distance on round spheres is also CND.
-In the construction developed in Section 3, however, conditional negative definiteness can be established directly from the structure of the proposed distance, substantially simplifying this verification step for set-indexed domains.
+This implication allows covariance constructions developed in Euclidean settings to be transferred directly to a wide range of non-Euclidean domains encountered in spatial statistics, including those based on $\alpha$-norms (Richards,, 1985), great-circle distance (Meckes,, 2013) and graph-indexed spaces (Anderes et al.,, 2020).
 3 Isotropic positive definite functions for sets
-To motivate the proposed distance, we first consider the classical Hausdorff distance, which provides a natural notion of discrepancy between spatial sets but lacks general guarantees for valid covariance construction.
+We address the general problem of constructing valid isotropic dependence structures on collections of spatial sets, rather than on point-indexed domains.
+As a natural starting point, one may consider the classical Hausdorff distance, which provides a natural notion of discrepancy between spatial sets but lacks general guarantees for valid covariance construction.
 For non-empty sets $A_{1},A_{2}\subset D$, the Hausdorff distance is defined as
@@
 While $d_{h}$ constitutes a valid metric on the class of non-empty, closed, and bounded subsets of $D$ (Arutyunov et al.,, 2016), it presents significant challenges for statistical modeling.
-Despite its intuitive interpretation in spatial statistics (Godoy et al.,, 2022, 2025), computing the Hausdorff distance is often prohibitively expensive (Knauer et al.,, 2011; Parikh,, 2023).
-More critically, to the best of our knowledge, constructing valid positive-definite functions based directly on $d_{h}$ remains an open problem.
+Despite its intuitive interpretation in spatial statistics (Godoy et al.,, 2022; Godoy et al., 2026a, ), computing the Hausdorff distance is often prohibitively expensive (Knauer et al.,, 2011) and, to the best of our knowledge, constructing valid positive-definite functions based directly on $d_{h}$ remains an open problem.
 These analytical and computational limitations motivate the ball-Hausdorff distance defined below.
-3.1 The ball-Hausdorff distance
 We define the ball–Hausdorff distance by reducing each spatial set to its minimal enclosing ball, thereby retaining only location and scale information relevant for dependence modeling.
@@
 For the unit sphere, this requires a Chebyshev radius less than $\pi/2$ (Bridson and Haefliger,, 1999, Proposition 1.4).
-This condition satisfied by all spatial configurations in our numerical experiments.
-Using the definition 3.1, we define the geometric summary $\mathcal{B}(A)$ as the minimum enclosing ball of $A$, characterized by the center $c(A)$ and radius $R(A)$.
+This condition is satisfied by all spatial configurations in our numerical experiments.
+Using Definition 3.1, we define the geometric summary $\mathcal{B}(A)$ as the minimum enclosing ball of $A$, characterized by the center $c(A)$ and radius $R(A)$.
 Note that, if $A$ is already a ball (or a singleton), it follows trivially that $A=\mathcal{B}(A)$.
@@
 The ball–Hausdorff distance inherits basic pseudometric properties from the Hausdorff distance while remaining insensitive to within-set geometry.
-In particular, for $A_{1},A_{2},A_{3}\in\mathcal{C}_{D}$,
-$bh(A_{1},A_{2})$ $\geq 0,$
-$bh(A_{1},A_{2})$ $=0,\quad\text{ if }\quad A_{1}=A_{2},$
-$bh(A_{1},A_{2})$ $\leq bh(A_{1},A_{3})+bh(A_{2},A_{3}),$
-where the first two properties are immediate from the definition and the third follows from the fact that Hausdorff distance is a metric on $\mathcal{C}_{D}$ (Molchanov,, 2005, Appendix C).
+In particular, non-negativeness follows immediately from the definition, while the triangle inequality is a consequence of the fact that Hausdorff distance is a metric on $\mathcal{C}_{D}$ (Molchanov,, 2005, Appendix C).
 Note that the distance does not render a metric because $bh(A_{1},A_{2})=0$ does not imply $A_{1}$ and $A_{2}$ are identical.
@@
 In other words, the proposed distance function between points reduces to $d$, from the original metric space.
-In Appendix C, we also derive an upper-bound for the ball-Hausdorff distance.
+In Appendix B, we also derive an upper-bound for the ball-Hausdorff distance.
 An important question in practice is how to compute the ball-Hausdorff distance.
@@
 For instance, calculating the Hausdorff distance between two balls may not always be trivial (see Section 2 for the problems in calculating the Hausdorff distance).
-In the special case where the study region lies on the Euclidean plane, a closed-form formula for the Hausdorff distance between circles has been developed (Scitovski and Marošević,, 2015).
-We extend this formula to length spaces which is general enough to cover most scenarios relevant for spatial statistics applications.
+In the special case where the study region lies on $\mathbb{R}^{2}$, a closed-form formula for the Hausdorff distance between circles has been developed (Scitovski and Marošević,, 2015).
+We extend this formula to length spaces, which are general enough to cover most scenarios relevant for spatial statistics applications.
 Theorem 3.3.
@@
 In the limiting case where sets have identical radii or are small relative to their separation, the distance reduces to the distance between their centers.
-3.2 Isotropic positive definite functions for sets
+Despite their structural differences, the ball–Hausdorff and Hausdorff distance may coincide.
+The corollary below presents one such instance.
+Corollary 3.3.1.
+In $\mathbb{R}^{1}$, $bh(\cdot,\cdot)\equiv d_{h}(\cdot,\cdot)$.
+Proof.
+Let $[a,b]$ and $[c,d]$ represent two arbitrary intervals in $\mathbb{R}^{1}$. Then, we have:
+$bh([a,b],[c,d])$ $=\frac{1}{2}[\lvert(a+b)-(c+d)\rvert+\lvert(b-a)-(d-c)\rvert$]
+$=\frac{1}{2}[\lvert(b-d)+(a-c)\rvert+\lvert(b-d)-(a-c)\rvert$]
+[…]
+Similarly,
+$d_{h}([a,b],[c,d])$ $=\inf\{r\geq 0\,:\,[a,b]\subseteq B_{r}([c,d]),[c,d]\subseteq B_{r}([a,b])\}$
+$=\inf\{r\geq 0\,:\,[a,b]\subseteq[c-r,d+r],[c,d]\subseteq[a-r,b+r]\}$
+[…]
+Therefore,
+$bh([a,b],[c,d])=d_{h}([a,b],[c,d]),$
+as claimed. ∎
 In addition to providing a simplified expression for the ball-Hausdorff distance, Theorem 3.3 enables a direct characterization of its conditional negative definiteness, which underpins our main result.
@@
 For interpretability, dependence is summarized using the practical range $\rho$, defined as the distance where correlation decays to 0.1.
-For the PEXP family, $\rho={\log(10)}^{1/\nu}\phi$ (Godoy et al.,, 2025), while for the Matérn family, $\rho\approx\phi\sqrt{8\nu}$ (Lindgren et al.,, 2011).
+For the PEXP family, $\rho={\log(10)}^{1/\nu}\phi$ (Godoy et al., 2026a, ), while for the Matérn family, $\rho\approx\phi\sqrt{8\nu}$ (Lindgren et al.,, 2011).
 These models form the basis for distance comparisons and eigenvalue-based diagnostics that follow.
@@
 Distances were rescaled to the unit interval by dividing by the maximum observed distance within each scenario, so that comparisons reflect relative rather than absolute scale differences across spatial configurations.
-Under this joint normalization, the values for of either metric need not span the full unit interval, as the observed minimum and maximum can be bounded away from 0 and 1 by geometric constraints inherent to the synthetic designs, such as shared boundaries, partial overlap, and limited variability in set geometry.
 Figure 2 displays two-dimensional kernel density estimates comparing the resulting distance values.
@@
 For the Matérn family, covariance matrices are constructed using square-root transformed distances, as required by Theorem 2.3.
-The practical range parameter is varied over $\rho\in[0.8,1.5]$, where the difference between the positive definiteness of the resulting covariance matrices from the two distances are notable.
-The smallest eigenvalue of each resulting covariance matrix, denoted $\lambda_{1}$, is used to assess positive definiteness, with negative values indicating invalid covariance structures.
+The practical range parameter is varied over $\rho\in[0.8,1.5]$, where the differences between the positive definiteness of the resulting covariance matrices from the two distances are notable.
+The smallest eigenvalue of each resulting covariance matrix, denoted $\lambda_{1}$, is used to assess lack of positive definiteness, with negative values indicating invalid covariance structures.
 Figure 3: Eigenvalue diagnostics for covariance matrices constructed using the PEXP family with $\nu=1$ (top row) and the Matérn family with $\nu=2.5$ (bottom row).
@@
 In all cases considered, covariance matrices based on the ball-Hausdorff distance remain positive definite, with $\lambda_{1}\geq 0$ throughout the examined range of $\rho$.
-In contrast, covariance matrices constructed using the standard Hausdorff distance frequently exhibit negative eigenvalues, even when square-root transformation is applied for the Matérn family, a reported limitations of Hausdorff-based constructions in set-indexed settings (Godoy et al.,, 2025).
+In contrast, covariance matrices constructed using the standard Hausdorff distance frequently exhibit negative eigenvalues, even when square-root transformation is applied for the Matérn family, a reported limitation of Hausdorff-based constructions in set-indexed settings (Godoy et al., 2026a, ).
 These violations are most pronounced for fused data and for configurations defined on the sphere.
@@
 These properties allow direct construction of valid set-indexed Gaussian processes for problems such as spatial data fusion and change-of-support, while still accommodating stochastic integrals when scientifically appropriate.
-An interesting theoretical direction is the development of valid cross-covariance functions for sets (Genton and Kleiber,, 2015), which would support modeling when responses and explanatory variables are observed on different spatial supports.
 Several theoretical questions remain open.
@@
 32), but extending this notion to set-indexed or manifold-indexed processes requires more general concepts of differentiation (Khastan et al.,, 2021).
-This issue is non-trivial, even for Gaussian processes on spheres (Alegría et al.,, 2021; Guinness and Fuentes,, 2016), and while recent work has defined mean-square differentiability for Gaussian processes on manifolds using smooth paths (Lin et al.,, 2019), a systematic treatment in the present setting remains to be developed.
+This issue is non-trivial, even for Gaussian processes on spheres (Alegría et al.,, 2021), and while recent work has defined mean-square differentiability for Gaussian processes on manifolds using smooth paths (Lin et al.,, 2019), a systematic treatment in the present setting remains to be developed.
 A further limitation arises from the pseudometric nature of the ball–Hausdorff distance: distinct sets with identical centers and radii are indistinguishable, which may lead to singular covariance matrices in degenerate cases.
@@
 Prates acknowledges the research grants obtained from CNPq-Brazil (309186/2021-8), FAPEMIG (APQ-01837-22, APQ-01748-24), and CAPES, respectively, for partial financial support.
-Appendix A Proofs
-A.1 Lemma 2.1
-Proof.
-Let $u\in\mathrm{B}_{r}(\mathrm{B}_{k}(x))$. Then $d(u,v)<r$ for some $v\in\mathrm{B}_{k}(x)$, so $d(x,v)<k$. By the triangle inequality,
-$d(x,u)\leq d(x,v)+d(v,u)<k+r,$
-thus, $u\in\mathrm{B}_{r+k}(x)$.
-Conversely, let $u\in\mathrm{B}_{r+k}(x)$.
-Note that $(D,d)$ is a length space (by assumption).
-In length spaces shortest paths always exist (Burago et al.,, 2001, Lemma 2.4.8).
-Thus, let $\gamma$ be a shortest path from $x$ to $u$.
-Let $v$ be the point on $\gamma$ at distance $k$ from $x$.
-Then $d(x,v)=k$ and d(u,v)=d(x,u)k<rd(u,v)=d(x,u)-k<r.
-Therefore, $v\in\mathrm{B}_{k}(x)$ and $u\in\mathrm{B}_{r}(v)$, proving $\mathrm{B}_{r+k}(x)\subseteq\mathrm{B}_{r}(\mathrm{B}_{k}(x))$.
-∎
-A.2 Corollary 2.2.1
-Proof.
-If $d$ is CND, Theorem 2.2 ensures that $(D,d^{1/2})$ can be isometrically embedded in a Hilbert space $H$.
-Consequently, there exists a map $\phi:D\to\mathbb{R}^{p}$ such that $d(s,t)=\lVert\phi(s)-\phi(t)\rVert_{2}$ for all $s,t\in D$.
-By Corollary 4.8 of Wells and Williams, (1975), the function $\exp\{-\kappa\lVert\cdot\rVert_{2}^{2\nu}\}$ is RPD on $L_{2}$ for $\nu\in(0,1]$.
-Therefore, for any set of points $s_{1},\dots,s_{m}\in D$ and real numbers $a_{1},\dots,a_{m}$, we have
-$\sum_{i=1}^{m}\sum_{j=1}^{m}a_{i}a_{j}\exp\{-\kappa d(s_{i},s_{j})^{\nu}\}$ $=\sum_{i=1}^{m}\sum_{j=1}^{m}a_{i}a_{j}\exp\{-\kappa\lVert\phi(s_{i})-\phi(s_{j})\rVert_{2}^{2\nu}\}\geq 0.$
-This establishes that $\exp\{-\kappa d^{\nu}\}$ is RPD on $(D,d)$. ∎
-Appendix B Length and geodesic spaces
+Appendix A Length and geodesic spaces
 A particularly important class of metric spaces are length spaces, where the distance is determined by paths connecting points.
@@
 A length structure consists of a class of “admissible paths” (e.g., all continuous paths) for which we can measure a non-negative length.
-Definition B.1 (Definition 2.3.1 in Burago et al., (2001)).
+Definition A.1 (Definition 2.3.1 in Burago et al., (2001)).
 Let $(D,d)$ be a metric space and $\gamma$ a continuous map $\gamma\,:\,[a,b]\to D$.
@@
 $d_{L}(x,y)=\inf\{L_{d}(\gamma)\,;\,\gamma:[a,b]\to D,\,\gamma\in A,\gamma(a)=x,\gamma(b)=y\}.$
-When the construction above induces the initial metric $d$ (i.e., $d_{L}=d$), than it is said to be intrinsic (Burago et al.,, 2001).
+When the construction above induces the initial metric $d$ (i.e., $d_{L}=d$), then it is said to be intrinsic (Burago et al.,, 2001).
 Although our primary motivation for working with length spaces is theoretical, they are broad enough to encompass the practical statistical challenges we address.
@@
 It is also worth noting that shortest paths need not be unique.
-Appendix C An upper bound for the ball-Hausdorff distance
+Appendix B An upper bound for the ball-Hausdorff distance
 In this section, we utilize an alternative formulation of the Hausdorff distance. Specifically, contrasting with Equation (1), we express the metric as (Arutyunov et al.,, 2016; Barrera et al.,, 2021)
@@
 A closed-form upper bound for the ball-Hausdorff distance is available under the mild assumption that $d$ is a pseudometric.
-Lemma C.1.
+Lemma B.1.
 Suppose $(D,d)$ is a pseudometric space.